OpenStudy (anonymous):

How would you find the solution to the equation x^4-64=0?

OpenStudy (anonymous):

Diffrence of squares. Factor away!

OpenStudy (zzr0ck3r):

(a^2-b^2) = (a-b)(a+b) here a = x^2 and b = 8 so (x^2-8)*(x^2+8) make sense?

OpenStudy (anonymous):

add 64 to the other side. 4 root of both sides. so 4th root of 64 = x

OpenStudy (zzr0ck3r):

nono cant do that @crazykidguy

OpenStudy (zzr0ck3r):

we will have 4 answers

OpenStudy (zzr0ck3r):

2 with imaginary parts

OpenStudy (zzr0ck3r):

+-sqrt(-8) and +-sqrt(8)

OpenStudy (usukidoll):

Factor